# Live Video Captioning

### v1

@@ III-C Novel Evaluation Metric for Live Video Captioning: the Live Score
III-C1 Online evaluation metric properties
The main characteristics of traditional offline dense video captioning metrics are as follows. All the information from the video annotations is introduced at once in the metrics. Then, the scorer rates the entire input video regardless of its duration. This scenario is quite different from that of LVC, where dense caption predictions with their associated timestamps arrive in temporal blocks, whose duration is less than that of the full video being processed. Moreover, the offline metrics cast a score that corresponds to the average score for a full video. This fact does not allow us to observe a temporal evolution in the performance of the models. This point is interesting for LVC because if we analyze the metric results with a temporal evolution, we can allow it to recover or deteriorate over time. In other words, we need an online nature metric that evolves with the video and reflects the real-time and online accuracy of LVC systems.
@@ III-C1 Online evaluation metric properties
  • Video stream based scoring: Ability to obtain scores from a video stream, therefore no access to a whole video is needed.
@@
III-C2 The Live Score
For the new paradigm of LVC, we propose a purely online metric: the Live Score (LS). In short, it is an adaptation of the various scores used in offline metrics, but tailored to process video streams online, causally, and with history, as we have specified in the previous section.
@@ III-C2 The Live Score
The proposed LS metric has an online nature, meeting all the requirements detailed in Section III-C1. However, it does not take into account the influence of false positives, i.e., predictions of captions that do not appear in the annotations available in the database. In other words, the proposed metric does not include any calibration mechanism with respect to false positives. In a realistic scenario for the LVC problem, such as generating captions for a live video stream from a surveillance camera, it is highly likely that there are large portions of the video where no action is occurring. Thus, in an LVC model applied to a video surveillance system, we must avoid at all costs the model generating captions for events that have not occurred. An LVC system will be accurate if it provides accurate captions, but also if it only provides captions when something relevant is happening in the video. To calibrate our metric and make it sensitive to false positives, we propose integrating a penalty for false positives into the LS. This new version of the metric is called weighted-LS (wLS), and is formulated as follows:
@@
where we have added the correction factor $\beta$, dependent on $fp(t_{n}^{\prime})$, which is number false positives corresponding to the video segment associated to $t_{n}^{\prime}$.
The two new proposed metrics, LS and its calibrated version wLS, allow for the online and continuous evaluation of what happens in a video up to time $t^{\prime}$, taking into account the entire history of the video from when it started processing at $t=0$. It may happen that the evaluation process starts with predictions that have very low scores and gradually improve over time, or vice versa. In such scenarios, since the metric is calculated based on all previous scores obtained by the system from $t=0$, it always considers the entire temporal timeline, and the metric may fail to reflect the system’s most recent behavior. To address this issue, we propose a version for both LS and wLS that considers only a fixed temporal history window for computing the performance of the LVC system at time $t^{\prime}$, covering only the interval $[t^{\prime}-w\Delta t,t^{\prime}]$. Here, $w$ defines the size of this fixed temporal window used to compute the metrics LS and wLS. By updating the metric considering only the fixed temporal history window, we allow it to evolve, reflecting the current performance of the LVC model. In Figure 5, we illustrate the calculation process for a temporal window size of $w=5$.
@@
IV Experimental Evaluation
In this section, we provide details of the experimental evaluation designed for the proposed LVC problem.
We start in Section IV-A with a description of the experimental setup, where we outline the database used and the adaptations made to it for the evaluation of the live models. Subsequently, in Sections IV-B-IV-F, we include both qualitative and quantitative results of all our experiments. The questions we want to address with the proposed experimental evaluation are as follows:
@@ IV-A Experimental setup
The different $\Delta t$ values used in the experiments were $(24,48,72,96,120,150)$ in frame numbers. We justify this choice of values for $\Delta t$ because all of them represent a reasonable temporal length for real-time applications. If the values were greater, the delay between predictions would be too high, making them unsuitable for consideration as Live models.
IV-B Evaluating LVC models with off-line metrics
Are offline experimental evaluation metrics suitable for LVC models? This is the question we want to specifically address in this section. We set up an experimental evaluation scenario in which we use the official offline metrics provided in the ActivityNet Challenge 2018 [39], on the online predictions generated by our LVC system. For localization performance, the average precision, average recall across intersection over union at different thresholds are used. For dense captioning performance, the official evaluation tool provided by ActivityNet Challenge 2018 is followed, which calculates the average precision measured by BLEU4, METEOR, and ROUGE_L scorers, of the matched pairs between generated captions and the ground truth across intersection over union thresholds of 0.3, 0.5, 0.7 and 0.9.
In Table I, we first present the results in terms of dense captioning accuracy. We provide a detailed comparison between the results of our online LVC model and those offered by the offline models PDVC [10] and Vid2Seq [9], that can be considered as the representatives of the state of the art for the dense video captioning problem.
@@ IV-B Evaluating LVC models with off-line metrics
TABLE I: Experimental comparison in terms of accuracy of the captions generated by offline models and our online model LVC using traditional offline metrics for the ActivityNet Captions dataset, using the validation set of videos.
@@
We can also compare offline and live models in terms of the accuracy of temporal localization of the captions, again using traditional offline metrics: precision and recall. In Table II, we present this detailed analysis, and we can observe that the results of event localization obtained for our predictions using these offline metrics are not satisfactory. Again, the metrics improve as $\Delta t$ increases in our LVC model. The explanation is similar to what we have provided for the previous metrics: our caption predictions are associated with video segments of duration $\Delta t$, which causes them not to meet the intersection-over-union criteria employed by the offline metrics. Observe the low performance when the threshold of 0.9 is used for the intersection-over-union.
@@
TABLE II: Caption localization for the validation video set of ActivityNet Captions, using traditional offline metrics. Comparison with the state of the art offline models.
IV-C Analysis using modified offline annotations
One could argue that offline metrics could still be used in an online scenario if annotations in the videos are modified so that their duration matches that used by live models. We have also performed this analysis for completeness, although we anticipate that this approach has significant drawbacks.
@@ IV-C Analysis using modified offline annotations
We have automated a process to modify the annotations provided in the validation set of the ActivityNet Captions database. Once the modified annotations are generated, traditional metrics for offline video captioning are employed, and the results are as follows. In Table III, we present the results in terms of caption localization in this new scenario and compare it with the performance obtained with the original annotations. Note that by splitting the provided annotations, we ensure that the predictions of our live model are not filtered out because they do not meet the intersection over union criterion.
@@
TABLE III: Comparison of results using original annotations or annotations adapted to live models, in terms of caption localization.
We also report the performance of our model for the three different metrics used to evaluate the precision of the captions, see Table IV. Again, we compare the performance when the original annotations and the adapted ones are used. We can observe that in this experiment the trend is that, the lower $\Delta t$ is, the higher the results are.
@@
TABLE IV: Comparison between results using original annotations and annotations adapted to the live scenario for ActivityNet Captions, employing LVC as the model.
In conclusion, adapting annotations to how live models work has a positive impact on their evaluation. However, the proposed adaptation has the drawback that the annotation must be dynamically adjusted to the temporal window $\Delta t$ being used, to then employ traditional offline metrics. Furthermore, it is not an evaluation strategy that naturally provides a metric that allows us to observe the temporal evolution of the model. In other words, these offline metrics do not evolve over time with the video, a fundamental aspect for the novel LVC problem. All these drawbacks are clearly addressed by the new metric proposed in this work.
@@
IV-D Evaluating LVC models with the online metric: the LiveScore
One of the main motivations of our work has been to design a new online evaluation metric for live video captioning models, i.e., the LiveScore (LS) (see Section III-C2). In this section, we detail all the experimental evaluation carried out using it. Note that we provide a detailed experimental analysis considering the four alternatives proposed for the LS metric:
@@ IV-D Evaluating LVC models with the online metric: the LiveScore
  • Mode with correction factor (Weighted Live Score - wLS).
  • Mode with history in memory window (LS with History Window - hLS).
  • Combined mode with correction factor and history in memory window (Weighted LS with History Window - whLS).
In Table V, we show the mean obtained by LS when integrated with the different scorers Bleu4, METEOR, and ROUGE_L. Based on the results obtained, we can draw the following conclusions. First, it is observed that the metrics increase as the parameter $\Delta t$ increases. This makes sense, as the larger the temporal window, the LVC model is able to see more portions of the video and offer a better description in the caption. However, values higher than $\Delta t=120$ do not seem to offer a significant improvement. Second, the results reported by the online metric are considerably higher than those obtained by the offline metrics. This becomes evident when comparing the results between Table I and Table V, where the results with the offline and online metrics are shown, respectively. For example, for $\Delta t=120$, the Bleu4 scorer goes from $1.55$ to $19.02$ for the LS version, or to $18.19$ for the wLS version. The increase experienced for the rest of the scorers is somewhat similar. This demonstrates the suitability of the new metric for the online scenario. Third, when comparing between the LS and wLS versions, we observe how LS offers higher results than the wLS version. This is because false positives that are generated are not considered by LS, but only by the wLS version, thus offering lower scores, but more adjusted to the actual performance of the LVC model. Fourth, in view of the results offered by the different scorers and the LS and wLS metrics, the LVC model achieves the best results for $\Delta t=[72,120]$. If we consider wLS as a more reliable metric, in the sense that it incorporates the penalty for false positives, LVC offers its best performance for $\Delta t=120$, with consensus among all scorers.
@@
LS Bleu4 18.03 18.79 18.99 19.01 19.02 18.75
METEOR 8.93 9.22 9.24 9.22 9.21 9.04
ROUGE_L 19.73 20.57 20.81 20.80 20.80 20.62
wLS Bleu4 16.53 17.58 17.94 18.06 18.19 18.02
METEOR 8.19 8.61 8.71 8.75 8.80 8.68
ROUGE_L 18.13 19.22 19.62 19.74 19.87 19.80
TABLE V: Evaluation for the LVC model using the new online metric. LS: Live Score. wLS: Weighted LS. We evaluate the LVC model for different values of the parameter $\Delta t$, as well as integrating different scorers (Bleu4, METEOR, and ROUGE_L).
Finally, we include the analysis using the hLS and whLS metrics. It should be noted that these were designed to be able to continuously visualize the accuracy of live video captioning systems. Therefore, they allow us to generate graphs where we can follow the temporal evolution of the different scorers. In Figures 6 to 11, we show the results obtained for 3 videos and all the values of the parameter $\Delta t$ used. Analyzing these graphs, we can draw the following conclusions. Firstly, the granularity of the metrics naturally increases as we decrease the parameter $\Delta t$. This allows us to control the speed at which captions are generated and the speed at which their accuracy is evaluated. Secondly, we have also included LS and wLS metrics without the temporal window option in these graphs. We can observe how for all videos and all scorers, the wLS version is always more conservative, reporting lower or equal scores, as it applies a penalty based on false positives. The same behavior is observed when comparing hLS and whLS, with the latter offering the most conservative scores. It is also noticeable that all versions start by reporting exactly the same values, beginning to diverge when the history window comes into play. As the third conclusion, perhaps the most important one, these plots show how the inclusion of a history window in our metric allows the model to exhibit its recovery in terms of caption generation accuracy. It no longer considers the entire past, only the recent local past, so that the scorers can increase or recover as the LVC model chains more and more correct caption predictions. In fact, observing a downward trend in the hLS and whLS metrics is an important indicator of how poorly the caption generator system is performing.
@@
IV-E Qualitative results
In this section, we present some qualitative results of our LVC system. Specifically, we show two examples corresponding to two videos from the validation set: one demonstrating a good result and the other a poor result. We used $\Delta t=150$, the maximum value we experimented with. This makes it easier to identify points in the graph where there is a change in slope according to the proposed metrics. We demonstrate the operation of the metric using the Live Score strategy, integrating the Bleu4 metric. In both examples, we mark several points on the graph where a change in slope is observed. If we observe a pronounced change in slope in the graph, it is because the score at that point improves or worsens considerably compared to previous points. Thus, the points marked on the graph determine points of interest, as they represent where the predictions resemble the ground truth more or less. The red shading indicates that the prediction is less similar, while the green shading indicates the opposite.

### v2

@@ 3.3 Novel Evaluation Metric for Live Video Captioning: the Live Score
3.3.1 Online evaluation metric properties
The main characteristics of traditional offline dense video captioning metrics are as follows. All the information from the video annotations is introduced at once in the metrics. Then, the scorer rates the entire input video regardless of its duration. This scenario is quite different from that of LVC, where dense caption predictions with their associated timestamps arrive in temporal blocks, whose duration is less than that of the full video being processed. Moreover, the offline metrics cast a score that corresponds to the average score for a full video. This fact does not allow us to observe the temporal evolution in the performance of the models, and this property is interesting for LVC. In other words, we need an online nature metric that evolves with the video and reflects the online accuracy of LVC systems.
@@ 3.3.1 Online evaluation metric properties
  • Video stream-based scoring: Ability to obtain scores from a video stream, therefore no access to a whole video is needed.
@@
3.3.2 The Live Score
For the new paradigm of LVC, we propose a purely online metric: the Live Score (LS). In short, it is an adaptation of the various scores used in offline metrics, but tailored to process video streams online, causally, and considering the history of the video, as we have specified in the previous section. Evaluating captioning quality in live video scenarios requires continuous assessment, as the accuracy of captions must be ensured at each stage of video playback. We propose an experimental evaluation guided this novel LS metric, which tracks the temporal evolution of caption accuracy as the video progresses. The LS metric allows us to evaluate caption performance at each moment, taking into account both the immediate accuracy and the cumulative quality of predictions. This continuous evaluation is crucial in live scenarios, where the model must maintain high performance throughout the video’s duration, ensuring that captions remain relevant and precise as events unfold.
@@ 3.3.2 The Live Score
The proposed LS metric has an online nature, meeting all the requirements detailed in Section 3.3.1. However, it does not take into account the influence of false positives, i.e., predictions of captions that do not appear in the annotations available in the database. In other words, the proposed metric does not include any calibration mechanism with respect to false positives. In a realistic scenario for the LVC problem, such as generating captions for a live video stream from a surveillance camera, it is highly likely that there are large portions of the video where no action is occurring. Thus, in an LVC model applied to a video surveillance system, we must avoid at all costs the model generating captions for events that have not occurred. Indeed, false-positive captions pose a significant challenge in live video captioning where they can disrupt the user experience by generating frequent, inaccurate descriptions. An LVC system will be accurate if it provides accurate captions, but also if it only provides captions when something relevant is happening in the video.
To calibrate our metric and make it sensitive to false positives, we propose integrating a penalty for false positives into the LS. This new version of the metric is called weighted-LS (wLS), and is formulated as follows:
@@
where we have added the correction factor $\beta$, dependent on $fp(t_{n}^{\prime})$, which is the number false positives corresponding to the video segment associated to $t_{n}^{\prime}$. By incorporating this weighted metric, our experimental evaluation pipeline enhances the reliability of live captioning systems, ensuring a more seamless and coherent user experience.
The two new proposed metrics, LS and its calibrated version wLS, allow for the online and continuous evaluation of what happens in a video up to time $t^{\prime}$, considering the entire history of the video from $t=0$. It may happen that the evaluation process starts with predictions that have very low scores and gradually improve over time, or vice versa. In such scenarios, since the metric is calculated based on all previous scores obtained by the system from $t=0$, it always considers the entire temporal timeline, and the metric may fail to reflect the system’s most recent behavior. To address this issue, we propose a version for both LS and wLS that considers only a fixed temporal history window for computing the performance of the LVC system at time $t^{\prime}$, covering only the interval $[t^{\prime}-w\Delta t,t^{\prime}]$. Here, $w$ defines the size of this fixed temporal window used to compute the metrics LS and wLS. By updating the metric considering only the fixed temporal history window, we allow it to evolve, reflecting the current performance of the LVC model. In Figure 5, we illustrate the calculation process for a temporal window size of $w=5$.
@@
4 Experimental Evaluation
In this section, we provide details of the experimental evaluation designed for the proposed novel LVC problem.
We start in Section 4.1 with a description of the experimental setup, where we outline the database used and the adaptations made to it for the evaluation of the live models. Subsequently, in Sections 4.2-4.6, we include both qualitative and quantitative results of all our experiments. The questions we want to address with the proposed experimental evaluation are as follows:
@@ 4.1 Experimental setup
The different $\Delta t$ values used in the experiments were $(24,48,72,96,120,150)$ in frame numbers. We justify this choice of values for $\Delta t$ because all of them represent a reasonable temporal length for online applications. If the values were greater, the delay between predictions would be too high, making them unsuitable for consideration as live models.
4.2 Evaluating LVC models with off-line metrics
Are offline experimental evaluation metrics suitable for the novel LVC problem? This is the question we want to specifically address in this section. We set up an experimental evaluation scenario in which we use the official offline metrics provided in the ActivityNet Challenge 2018 [41], on the online predictions generated by our LVC system. For localization performance, the average precision and average recall across intersection over union at different thresholds are used. For dense captioning performance, the official evaluation tool provided by ActivityNet Challenge 2018 is followed, which calculates the average precision measured by BLEU4, METEOR, and ROUGE_L scorers, of the matched pairs between generated captions and the ground truth across intersection over union thresholds of 0.3, 0.5, 0.7 and 0.9.
In Table 2, we first present the results in terms of dense captioning accuracy. We provide a detailed comparison between the results of our online LVC model and those offered by the offline models PDVC [10] and Vid2Seq [9], that are the state of the art for the dense video captioning problem.
@@ 4.2 Evaluating LVC models with off-line metrics
Table 2: Experimental comparison in terms of accuracy of the captions generated by offline models and our online model LVC using traditional offline metrics for the ActivityNet Captions dataset, using the validation set of videos.
@@
We can also compare offline and live models in terms of the accuracy of temporal localization of the captions, again using traditional offline metrics: precision and recall. In Table 3, we present this detailed analysis, and we can observe that the results of event localization obtained for our predictions using these offline metrics are not satisfactory. Again, the metrics improve as $\Delta t$ increases in our LVC model. The explanation is similar to what we have provided for the previous metrics: our caption predictions are associated with video segments of duration $\Delta t$, which causes them not to meet the intersection-over-union criteria employed by the offline metrics. Observe the low performance when the threshold of 0.9 is used for the intersection-over-union.
@@
Table 3: Caption localization for the validation video set of ActivityNet Captions, using traditional offline metrics. Comparison with the state of the art offline models.
4.3 Analysis using modified offline annotations
One might argue that offline metrics could still be applicable in an online scenario if the video annotations are adjusted to align with the duration used by LVC models. We have also performed this analysis for completeness, although we anticipate that this approach has significant drawbacks.
@@ 4.3 Analysis using modified offline annotations
We have automated a process to modify the annotations provided in the validation set of the ActivityNet Captions database. Once the modified annotations are generated, traditional metrics for offline video captioning are employed, and the results are as follows. In Table 4, we present the results in terms of caption localization in this new scenario and compare it with the performance obtained with the original annotations. Note that by splitting the provided annotations, we ensure that the predictions of our live model are not filtered out because they do not meet the intersection over union criterion.
@@
Table 4: Comparison of results using original annotations or annotations adapted to live models, in terms of caption localization.
We also report the performance of our model for the three different metrics used to evaluate the precision of the captions, see Table 5. Again, we compare the performance when the original annotations and the adapted ones are used. We can observe that in this experiment the trend is that, the lower $\Delta t$ is, the higher the results are.
@@
Table 5: Comparison between results using original annotations and annotations adapted to the live scenario for ActivityNet Captions, employing LVC as the model.
In conclusion, adapting annotations to how live models work has a positive impact on their evaluation. However, the proposed adaptation has the drawback that the annotation must be dynamically adjusted to the temporal window $\Delta t$ being used, to then employ traditional offline metrics. Furthermore, it is not an evaluation strategy that naturally provides a metric that allows us to observe the temporal evolution of the model. In other words, these offline metrics do not evolve over time with the video, a fundamental aspect for the novel LVC problem. All these drawbacks are clearly addressed by the new metrics proposed in this work.
@@
4.4 Evaluating LVC models with the novel Live Score
One of the main motivations of our work has been to design a new online evaluation metric for live video captioning models, i.e., the LiveScore (LS) (see Section 3.3.2). In this section, we detail all the experimental evaluations carried out using it. Note that we provide a detailed experimental analysis considering the four alternatives proposed for the LS metric:
@@ 4.4 Evaluating LVC models with the novel Live Score
  • Mode with a correction factor (Weighted Live Score - wLS).
  • Mode with history in a memory window (LS with History Window - hLS).
  • Combined mode with correction factor and history in a memory window (Weighted LS with History Window - whLS).
In Table 6, we show the mean obtained by the novel LS when integrated with the different scorers Bleu4, METEOR, and ROUGE_L. Based on the results obtained, we can draw the following conclusions. First, it is observed that the metrics increase as the parameter $\Delta t$ increases. This makes sense, as the larger the temporal window, the LVC model is able to see more portions of the video and offer a better description in the caption. However, values higher than $\Delta t=120$ do not seem to offer a significant improvement. Second, the results reported by the online metric are considerably higher than those obtained by the offline metrics. This becomes clear when comparing the results between Table 2 and Table 6, where the results with the offline and online metrics are shown, respectively. For example, for $\Delta t=120$, the Bleu4 scorer goes from $1.55$ to $19.45$ for the LS version, or to $18.96$ for the wLS version. The increase experienced for the rest of the scorers is somewhat similar. This demonstrates the suitability of the new metric for the online scenario. Third, when comparing between the LS and wLS versions, we observe how LS offers higher results than the wLS version. This is because false positives that are generated are not considered by LS, but only by the wLS version, thus offering lower scores, but more adjusted to the actual performance of the LVC model. Fourth, given the results offered by the different scorers and the LS and wLS metrics, the LVC model achieves the best results for $\Delta t=120$, with consensus among all scorers.
@@
LS Bleu4 18.13 19.00 19.29 19.38 19.45 19.25
METEOR 8.98 9.32 9.37 9.38 9.40 9.25
ROUGE_L 19.84 20.81 21.14 21.21 21.29 21.18
wLS Bleu4 17.12 18.19 18.61 18.80 18.96 18.84
METEOR 8.45 8.91 9.03 9.09 9.16 9.05
ROUGE_L 18.71 19.90 20.36 20.55 20.74 20.72
Table 6: Evaluation for the LVC model using the new online metric. LS: Live Score. wLS: Weighted LS. We evaluate the LVC model for different values of the parameter $\Delta t$, as well as integrating different scorers (Bleu4, METEOR, and ROUGE_L).
Finally, we include the analysis of the temporal evolution of the performance of our LVC model using all our novel evaluation metrics. This analysis demonstrates that the new metrics enable us to generate graphs that track the temporal evolution of different scores. Such tracking is crucial for LVC models and cannot be achieved with traditional offline metrics. Note that we include the hLS and whLS variations of the metric in this analysis, which are specifically designed to continuously visualize the accuracy of LVC systems.
In Figures 6 to 11, we show the results obtained for 3 videos and all the values of the parameter $\Delta t$ used. Analyzing these graphs, we can draw the following conclusions. Firstly, the granularity of the metrics naturally increases as we decrease the parameter $\Delta t$. This allows us to control the speed at which captions are generated and the speed at which their accuracy is evaluated. Secondly, we have also included LS and wLS metrics without the temporal window option in these graphs. We can observe how for all videos and all scorers, the wLS version is always more conservative, reporting lower or equal scores, as it applies a penalty based on false positives. The same behavior is observed when comparing hLS and whLS, with the latter offering the most conservative scores. It is also noticeable that all versions start by reporting exactly the same values, beginning to diverge when the history window comes into play. As the third conclusion, perhaps the most important one, these plots show how the inclusion of a history window in our metric allows the model to exhibit its recovery in terms of caption generation accuracy. It no longer considers the entire past, only the recent local past, so that the scorers can increase or recover as the LVC model chains more and more correct caption predictions. In fact, observing a downward trend in the hLS and whLS metrics is an important indicator of how poorly the caption generator system is performing.
@@
4.4.1 Discussion
While novel LS metric provides a robust online evaluation framework for LVC systems, it is not used directly during their training, as it can be observed in previous experiments. LS calculates a cumulative score over time, capturing the real-time and historical accuracy of predictions, which contrasts with the frame-by-frame optimization process used during training of our LVC model. Training objectives typically rely on instantaneous feedback, that are easier to optimize with stochastic gradient descent. Exploring methods to incorporate LS principles into a training objective remains an intriguing avenue for future research, potentially bridging the gap between evaluation and model optimization.
Another interesting aspect to analyze in the context of LVC is the quality of the annotations in the training datasets. The robustness of video captioning models heavily depends on the quality and diversity of annotations in the training datasets. High-quality annotations ensure accurate temporal boundaries and rich semantic details, enabling the model to learn fine-grained associations between actions and descriptions. Conversely, noisy or inconsistent annotations can introduce ambiguity, reducing the model’s ability to capture temporal precision and action differentiation. Similarly, a lack of variety in annotated events or scenarios limits the model’s capacity to generalize to unseen contexts, particularly in real-world applications with diverse and dynamic content. These factors are especially critical in LVC, where models must operate with partial observations and strict timing constraints. Addressing these limitations requires the development of datasets with comprehensive, high-quality annotations tailored to the online scenario.
4.4.2 Ablation Studies
In this section, we show the impact of the parameters of our LVC model on the quality of the dense captions. Specifically, we analyze the influence of the following components of our model: the use of a deformable transformer vs. a standard transformer; the number of queries; and the event counter head. For these studies we have used the METEOR metric in conjunction with our LS. Figure 12 shows the results obtained. As can be observed, the use of a deformable transformer significantly improves the results. Regarding the number of queries used in our transformer, the impact is not as substantial, with $q=10$ being a fairly efficient compromise. Lastly, it is notable that not using the event counter head leads to a significant drop in the model’s performance.
In summary, regarding the architecture used in the proposed LVC model, the use of a deformable transformer is confirmed as the option that provides the best results. The influence of the deformable transformer’s specific parameters, such as the number of queries, does not seem to significantly affect the quality of the captions. We believe this is because our model, instead of processing the entire video, works with short segments of the video, due to the online nature of the problem posed by LVC. Finally, the use of the three simultaneous heads—one for captions, another for localization, and another for considering the estimation of the number of events—appears to deliver the best results.
Figure 12: We show the impact of the different parameters and components of our LVC model on the quality of the dense captions when using a standard transformer (Vanilla) vs. a DETR, different values for $q$, and when the event counter head is not activated. We report METEOR with the LS metric.
4.5 Qualitative analysis
In this section, we present some qualitative results of our LVC system. Specifically, we show two examples corresponding to two videos from the validation set: one demonstrating a good result and the other a poor result. We used $\Delta t=150$, the maximum value we experimented with. This makes it easier to identify points in the graph where there is a change in slope according to the proposed metrics. We demonstrate the operation of the metric using the Live Score strategy, integrating the Bleu4 metric. In both examples, we highlight several points on the graph where notable changes in slope occur. A pronounced change in slope indicates a significant improvement or deterioration in the score compared to previous points. These marked points of interest represent where the predictions either align more closely or diverge from the ground truth. Red shading signifies a greater discrepancy between the prediction and the ground truth, while green shading indicates a closer match.